THEORETICAL EVIDENCE FOR THE FORMATION OF ROTATIONAL ENERGY LEVEL CLUSTERS IN THE VIBRATIONAL GROUND STATE OF PH_{3}

SERGEI N. YURCHENKO, WALTER THIEL, Max-Planck-Institut für Kohlenforschung, Kaiser-Wilhelm-Platz 1, D–45470 Mülheim an der Ruhr, Germany; SERGUEI PATCHKOVSKII, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario, Canada K1A 0R6; \underline{PER JENSEN}, FB C – Theoretische Chemie, Bergische Universität, D–42097 Wuppertal, Germany.

We investigate theoretically the rotational dynamics of pyramidal XY_{3} molecules in highly excited rotational states. Towards this end we compute, by a variational method,^{a} the rotational energy levels in the vibrational ground state of PH_{3} for J ≤ 80 from a potential energy surface determined previously.^{b} At J ≥ 50 the calculated energy levels show a distinct cluster pattern. By monitoring the cluster formation we follow the various stages of the rotational dynamics. We analyze the wavefunctions for the cluster states and compute expectation values which show that the C_{3v} geometrical symmetry of PH_{3} is broken at high rotational excitation. The conclusions drawn from the quantum-mechanical calculations are confirmed by semi-classical theory, i.e., by an analysis of the stationary points on the rotational energy surface.

aS. N. Yurchenko, M. Carvajal, P. Jensen, H. Lin, J. Zheng, and W. Thiel, Mol. Phys. 103, 359-378 (2005) and references therein.

bS. N. Yurchenko, M. Carvajal, P. Jensen, F. Herregodts, and T. R. Huet, Chem. Phys. 290, 59-67 (2003).